The pKb values for the dibasic base B are pKb1=2.10 and pKb2=7.52.

Calculate the pH at each of the points in the titration of 50.0 mL of a 0.70 M B(aq) solution with 0.70 M HCl(aq).

What is the pH before addition of any HCl?

pH=

What is the pH after addition of 25.0 mL HCl?

pH=

What is the pH after addition of 50.0 mL HCl?

pH=

What is the pH after addition of 75.0 mL HCl?

pH=

What is the pH after addition of 100.0 mL HCl?

pH=
Answer #1

Solution : Given that - dibasic Bese titration 500ml of 0.701 pkb =2.10 Ebi = 0.0079 piba = 7.52 kb = 1.39x108 Beau) Solution(b) pH affor addition of 2500 ml Hel Milimdes of B- 5080-70 = 35 mmoles, Millimoles of HCl = 25X0.70 = 1745 mmoles. Half waypH = H-7.52 TpH = 6+48) (C)pt After addition of 100.0 ml of Hcl. It is gond Fuilvalence point. gond Feuilvance point only acika XXX - 0.233- 7.57810* -wa 0233 xxx = 117.6 407 a = 4.12 x10 4 [Hgct] = k12 x10 4 M pH = -log (Hola x104) pH = 3:38.